OpenStudy (mathhelplllllllz2):

HELP PLEASE WILL MEDAL AND FAN Find the missing lengths of the sides.

OpenStudy (mathhelplllllllz2):



OpenStudy (mathhelplllllllz2):

A- A= 9 in., b = 9 in B- a = 9 in., b = 9 two squared C - a = 9 3 squared D.) 18 in. B= 18 in

OpenStudy (mathhelplllllllz2):

@TheSmartOne

TheSmartOne (thesmartone):

Is that \(\Large 9 \sqrt{2}\) for the side opposite the 90 degree angle?

OpenStudy (mathhelplllllllz2):

Yes

OpenStudy (cutegirl1000):

it is 45* thats your answer

OpenStudy (mathhelplllllllz2):

No it isn't

OpenStudy (mathhelplllllllz2):

Tthe answer choices are above

OpenStudy (mathhelplllllllz2):

@TheSmartOne that is \[9\sqrt{2}\]

TheSmartOne (thesmartone):

Use this: